Tullahoma Company purchased equipment for $27,500. It depreciated the equipment over a fiveyear life by the double-declining-balance method until the end of the second year, at which time the asset was sold for $8,500. Calculate the gain or loss on the sale at the end of the second year.Tullahoma Company purchased equipment for $27,500. It depreciated the equipment over a fiveyear life by the double-declining-balance method until the end of the second year, at which time the asset was sold for $8,500. Calculate the gain or loss on the sale at the end of the second year.


1
Expert's answer
2021-05-21T11:42:19-0400

Double declining rate

"\frac{100}{5}\times2
= 40%" %

Year 1

"\frac{40}{100}\times 27500
=11000"

=$11000

Ending carrying value

$27500 - $11000

= $16500

Year 2

"\frac{40}{100}\times 16500
= 6600"

= $ 6600

Ending carrying value

$16500 - $6600

= $9900

Loss of the sale at the end of Year 2

$9900 - $8500

= $1400

Answer= $ 1400